Please help me solve this equation in simple steps:
give the values of A, B and C needed to write the equation's general quadratic form.
2/3 (x-4)(x+5)=1

Answer by ewatrrr(24785)

 
Hi,
2/3 (x-4)(x+5)=1
(x-4)(x+5)= 3/2
x^2 + x - 20 - 3/2 = 0
x^2 + x -43/2 = 0
2x^2 + 2x - 43 = 0